John adding on to his house the original room a square with side length of x with the addition Journal at 10 feet to the length and 13 feet to the way the area of the new room will be 270 square feet what were the original dimensions of the room
Answer by ankor@dixie-net.com(22740)

John adding on to his house.
the original room is a square, with side length of x.
He adds 10 feet to the length and 13 feet to width.
the area of the new room will be 270 square feet
what were the original dimensions of the room
:
the new dimensions area equation
(x+10)*(x+13) = 270
FOIL
x^2 + 13x + 10x + 130 = 270
x^2 + 23x + 130 - 270 = 0
a quadratic equation
x^2 + 23x - 140 = 0
You can use the quadratic formula; a=1; b=23; c=-140, but this will factor to:
(x-5)(x+28) = 0
the positive solution
x = 5 ft is the side of the original square room
:
:
:
Check this. Find the new room dimension; 15 by 18 and area
15 * 18 = 270 sq/ft